OpenStudy (anonymous):

Find the absolute value of the complex number -3 - 6i. a. 6.07 c. 6.7 b. 6.17 d. 6.71 How would I find the absolute Value of the complex number?

OpenStudy (anonymous):

i can give u an example idk hope u can get help from that http://www.mathwarehouse.com/algebra/complex-number/absolute-value-complex-number.php

OpenStudy (anonymous):

so accordingly u can do it like this

OpenStudy (anonymous):

Find the absolute value of the complex number -3 - 6i Ans: sqrt[9+36] = sqrt(45) = 3sqrt(5)

OpenStudy (anonymous):

\[\left| -3 -6i \right|\]

OpenStudy (anonymous):

I found it. Thanks guys. ^^

OpenStudy (anonymous):

:D root over of 45